OpenStudy (anonymous):

The function f(x) = 20(2)x represents the growth of a butterfly population every year in a remote swamp. Jan wants to manipulate the formula to an equivalent form that calculates five times a year, not just once a year. Which function is correct for Jan's purpose and what is the new growth rate?

OpenStudy (anonymous):

So I know that for a year her growth rate is the f(x)=20(2)^x

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

By f(x) = 20(2)x, do you mean\[f(x) = 20(2)^x?\]

OpenStudy (anonymous):

Yes sir!

OpenStudy (mathmale):

I'm uncertain what you mean by your phrase "calculates five times per year." What does that mean to you? Calculate ... what?

OpenStudy (anonymous):

I am assuming that the function given is representing the formual for the growth rate of once a year..so she wants one that calculates five years?

OpenStudy (mathmale):

Let's compromise and express this as it is normally expressed: "rate of annual growth."

OpenStudy (anonymous):

Okay! Would you like me to write out the answer choices just so you can catch a glimpse of what exactly their needs are?

OpenStudy (mathmale):

Yes, but first let's discuss this problem a bit further. If, instead of typing\[f(x) = 20(2)^x\] we were to type \[f(x) = 20(2)^{5x}\]

OpenStudy (mathmale):

... how do you think the behavior of the original function would be modified?

OpenStudy (anonymous):

f(x) = 4(1.15)x; growth rate is 5% f(x) = 4(1.15)5x; growth rate is 115% f(x) = 4(2)x; growth rate is 200% f(x) = 4(2)x, growth rate is 5% Here are the choices, the numbers/letters after the parenthesis are ^. I believe the behvaior of the function would be modified by increasing the outcome?

OpenStudy (mathmale):

The outcomes would be dramatically larger were we to replace that \[2^x\] by \[2^{5x}\]

OpenStudy (mathmale):

If you think about what we've learned so far, could you now make an educated guess regarding which of the four possible answer choices is the correct one?

OpenStudy (mathmale):

Note that the problem you've typed in begins with a '20,' whereas all of the answer choices begin with a '4'.

OpenStudy (anonymous):

I sort of believe it would be b. Just because we added the 5x.

OpenStudy (anonymous):

That's a really high growth rate though

OpenStudy (mathmale):

Actually, you/we inserted the multiplier '5'...we did not 'add' the '5x'. In this case I'm unfortunately not clear on how we progressed from 20(2)^x to 4(x)^x. We have to account for that in some way.

OpenStudy (anonymous):

I'm really not sure either. I'm confused as to how one can just manipulate a formula

OpenStudy (mathmale):

As much as I'd like to help further, I don't yet see what was done with that '20'. I do see that if we start with f(x) = 20(2)^x, and insert the multiplier 5, we get \[f(x)=20(32)^x\]

OpenStudy (mathmale):

Manipulating a formula is perfectly legal and appropriate. What's not clear to me, and apparently not clear to you, is the significance of inserting that multiplier, '5'.

OpenStudy (anonymous):

Is it quite possible that the 20 is representing the amount of butterflies already? And I know that it's legal and appropritate, I'm just confused extrememly on how they're manipulating this problem is what I meant.

OpenStudy (mathmale):

Yes, it's true that the '20' in the original problem statement represents the original number of butterflies. Note that if x=0, 2^x = 2^0 =1 =f(0)=initial value of f(x). Were we to insert that factor of 5, the resulting formula would look like \[20(2)^{5x}=20(32)^x\] which would grow a LOT faster.

OpenStudy (mathmale):

a lot faster than the original formula would, I mean.

OpenStudy (anonymous):

Right, so now I see how you get the 20(32)^x.

OpenStudy (mathmale):

Unfortunately, I share your uncertainty in regard to how to manipulate this last expression so that it resembles one of the four possible answers.

OpenStudy (mathmale):

All I can suggest, at this point, is that you double check to ensure you've copied the problem statement correctly.

OpenStudy (mathmale):

In your shoes, I'd move on to a new problem and return to this one only if you can and if you have time. Choose whichever answer you think is best, based upon our discussions. I see you have indeed copied the problem down correctly. Wish I could help you take this problem all the way to a satisfactory, definite conclusion

OpenStudy (anonymous):

This is sadly the only question I have left to answer. Thank you for attempting to help, i highly appreciate it. Do you want me to tag some others to see if they can help us out this slump?

OpenStudy (mathmale):

Yes, I think you should. I look forward to working with you again in the near future (if not tonight).

OpenStudy (anonymous):

@SolomonZelman Do you have any idea how to do this problem? And most likely next school year! This is my final assignment needed for this years math course. :)